SPECTRAL LINESHAPES OF IR AMMONIA TRANSITIONS VERSUS PRESSURE AND TEMPERATURE

\underline{G. BALDACCHINI}, F. D’AMATO, M. DE ROSA, F. PELAGALLI, ENEA, Centro di Frascati, 00044 Frascati, Italy; G. BUFFA and O. TARRINI, Dipartimento di Fisica, Università di Pisa, 56100 Pisa, Italy.